An inverted regular square pyramid has an altitude of 12m and with top
edge of 3m is fully–filled with water.

4. What is the volume of the water in m3?
A portion of water is to be removed so that water will be 7m deep.
5. What is the area at the surface of the water when it is 7m deep?
1
Expert's answer
2020-11-17T17:36:15-0500

4. Given "a=3\ m, H=12\ m"


"V=\dfrac{1}{3}a^2H""V=\dfrac{1}{3}(3\ m)^2(12\ m)=36\ m^3"

The volume of the water is 36 m3.


5.


"\dfrac{a}{H}=\dfrac{a_1}{H_1}=const"

Then


"a_1=a\cdot\dfrac{H_1}{H}"

"a_1=3\ m\cdot\dfrac{7\ m}{12\ m}=1.75 \ m"

"S_1=(a_1)^2=(1.75\ m)^2=3.0625\ m^2"

The area at the surface of the water when it is 7m deep is 3.0625 m2.